Rhys and Gareth are having a tug-of-war with 1.25 kg of climbing rope. If Rhys pulls with a force of 24 N and the rope accelerates away from him at 1.4 m/s2 , with what force is Gareth pulling the rope?



1) Find the force in which Gareth is pulling the rope?



2) What is Rhys' mass in this situation?



(how much does he weight in kg)



(how much does he weigh if he pulls with the force of 24 N and the rope accelerates from him at 1.4 meters per second squared m/s2
1
Expert's answer
2018-10-31T14:05:09-0400

1) The force in which Gareth is pulling the rope is 24 N


2) Rhys' mass is

F=(m_R+m_rope )·a

m_R=F/a-m_rope=24/1.4-1.25=15.9 kg